# Finite-Sample Analysis of Policy Evaluation for Robust Average Reward Reinforcement Learning

### v2

@@ F.3 Formal Statement of Theorem 7
$\mathbb{E}$ $\left[\left|g_{t+1}-g^{\pi}_{\mathcal{P}}\right|\right]\leq(1-\beta_{t})\left|g_{t}-g^{\pi}_{\mathcal{P}}\right|+\beta_{t}\left(4\mathbb{E}\left[{\|V_{T}-V^{*}\|}_{\infty}\right]+\mathbb{E}[\left|\nu_{t}\right|]\right)$
$\leq(1-\beta_{t})\left|g_{t}-g^{\pi}_{\mathcal{P}}\right|+\beta_{t}\left(4\mathbb{E}\left[{\|V_{T}-V^{*}\|}_{\infty}\right]+\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|\right)$ (148)
@@
$+\zeta_{T}\sum_{t=0}^{T-1}(\frac{1}{\zeta_{t+1}})\beta_{t}\left(4\mathbb{E}\left[{\|V_{T}-V^{*}\|}_{\infty}\right]+\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|\right)$
$=\zeta_{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\sum_{t=0}^{T-1}(\frac{\zeta_{T}}{\zeta_{t+1}})\beta_{t}\left(4\mathbb{E}\left[{\|V_{T}-V^{*}\|}_{\infty}\right]+\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|\right)$
$\leq\zeta_{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\sum_{t=0}^{T-1}\beta_{t}{\left(4\mathbb{E}\left[{\|V_{T}-V^{*}\|}_{\infty}\right]+\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|\right)}$
$=\zeta_{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\left(4\mathbb{E}\left[{\|V_{T}-V^{*}\|}_{\infty}\right]+\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|\right)\sum_{t=0}^{T-1}\beta_{t}$ (149)
@@
$\mathbb{E}\left[\left|g_{T}-g^{\pi}_{\mathcal{P}}\right|\right]\leq\frac{1}{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\left(8\mathbb{E}\left[{\|V_{T}-V^{*}\|}_{\infty}\right]+2\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|\right)\log T$ (150)
Theorem 11 (Formal version of Theorem 7).
Following all notations and assumptions in Theorem 10, then for contamination uncertainty sets,
@@
Where all variables are defined the same as in Theorem 10.
@@
By Theorem 10, taking square root on both side and utilizing the concavity of square root function, we have for contamination uncertainty sets,
@@
Regarding the bound for the absolute bias of $\hat{\sigma}_{\mathcal{P}^{a}_{s}}$, from Lemma 12, we have for contamination uncertainty,
@@
In addition, combining (97)-(98) with Lemma 13, we have for for TV distance uncertainty,
@@ F.4 Proof of Theorem 7
We use the result from Theorem 11, to set $\mathbb{E}\Bigl{[}\left|g_{T}-g^{\pi}_{\mathcal{P}}\right|\Bigr{]}\leq\epsilon$. For contamination uncertainty sets we set $T=\operatorname*{\mathcal{O}}\left(\frac{t^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$, resulting in $\operatorname*{\mathcal{O}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$ sample complexity. For TV and Wasserstein uncertainty set, we set $N_{\max}=\operatorname*{\mathcal{O}}\left(\log\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$ and $T=\operatorname*{\mathcal{O}}\left(\frac{t^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log^{3}\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$, combining with Theorem 3, this would result in $\operatorname*{\mathcal{O}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log^{4}\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$ sample complexity.
Appendix G Some Auxiliary Lemmas for the Proofs
Lemma 10 (Theorem IV in Berger and Wang (1992)).
@@ Appendix G Some Auxiliary Lemmas for the Proofs
Let $\mathcal{Q}$ be a bounded set of square matrix such that $\rho(Q)<\infty$ for all $Q\in\mathcal{Q}$ where $\rho(\cdot)$ denotes the spectral radius. Then we have
$\sup_{Q\in\mathcal{Q}}\rho(Q)=\lim_{k\rightarrow\infty}\sup_{Q_{i}\in\mathcal{Q}}\rho(Q_{k}\ldots Q_{1})^{\frac{1}{k}}$ (160)
Lemma 11 (Lemma 6 in Zhang et al. (2021)).
@@
Under the setup and notation in Appendix D.1.1, if assuming the noise has bounded variance of $\mathbb{E}[\,\|w^{t}\|_{\infty,\overline{E}}^{2}|\mathcal{F}^{t}]\leq A+B\|x^{t}-x^{*}\|_{\infty,\overline{E}}^{2}$, we have
@@
Lemma 13 (Ergodic case of Lemma 9 in Wang et al. (2022)).
For any average reward MDP with stationary policy $\pi$ and the mixing time defined as

### v3

@@ Appendix D Convergence for Robust TD
$\mathbb{E}$ $\left[\left|g_{t+1}-g^{\pi}_{\mathcal{P}}\right|\right]\leq(1-\beta_{t})\left|g_{t}-g^{\pi}_{\mathcal{P}}\right|+\beta_{t}\left(4\mathbb{E}\left[\left\lVert V_{T}-V^{*}\right\rVert_{\infty}\right]+\mathbb{E}[\left|\nu_{t}\right|]\right)$
$\leq(1-\beta_{t})\left|g_{t}-g^{\pi}_{\mathcal{P}}\right|+\beta_{t}\left(4\mathbb{E}\left[\left\lVert V_{T}-V^{*}\right\rVert_{\infty}\right]+\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|\right)$ (180)
@@
$+\zeta_{T}\sum_{t=0}^{T-1}(\frac{1}{\zeta_{t+1}})\beta_{t}\left(4\mathbb{E}\left[\left\lVert V_{T}-V^{*}\right\rVert_{\infty}\right]+\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|\right)$
$=\zeta_{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\sum_{t=0}^{T-1}(\frac{\zeta_{T}}{\zeta_{t+1}})\beta_{t}\left(4\mathbb{E}\left[\left\lVert V_{T}-V^{*}\right\rVert_{\infty}\right]+\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|\right)$
$\leq\zeta_{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\sum_{t=0}^{T-1}\beta_{t}{\left(4\mathbb{E}\left[\left\lVert V_{T}-V^{*}\right\rVert_{\infty}\right]+\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|\right)}$
$=\zeta_{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\left(4\mathbb{E}\left[\left\lVert V_{T}-V^{*}\right\rVert_{\infty}\right]+\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|\right)\sum_{t=0}^{T-1}\beta_{t}$ (181)
@@
$\mathbb{E}\left[\left|g_{T}-g^{\pi}_{\mathcal{P}}\right|\right]\leq\frac{1}{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\left(8\mathbb{E}\left[\left\lVert V_{T}-V^{*}\right\rVert_{\infty}\right]+2\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|\right)\log T$ (182)
Theorem D.3 (Formal version of Theorem 6.2).
Following all notations and assumptions in Theorem D.2, then for contamination uncertainty sets,
@@
Where all variables are defined the same as in Theorem D.2.
@@
By Theorem D.2, taking square root on both side and utilizing the concavity of square root function, we have for contamination uncertainty sets,
@@
Regarding the bound for the absolute bias of $\hat{\sigma}_{\mathcal{P}^{a}_{s}}$, from Lemma F.3, we have for contamination uncertainty,
@@
In addition, combining (129)-(130) with Lemma F.4, we have for for TV distance uncertainty,
@@
We use the result from Theorem D.3, to set $\mathbb{E}\Bigl[\left|g_{T}-g^{\pi}_{\mathcal{P}}\right|\Bigr]\leq\epsilon$. For contamination uncertainty sets we set $T=\operatorname*{\mathcal{O}}\left(\frac{t^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$, resulting in $\operatorname*{\mathcal{O}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$ sample complexity. For TV and Wasserstein uncertainty set, we set $N_{\max}=\operatorname*{\mathcal{O}}\left(\log\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$ and $T=\operatorname*{\mathcal{O}}\left(\frac{t^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log^{3}\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$, combining with Theorem 5.1, this would result in $\operatorname*{\mathcal{O}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log^{4}\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$ sample complexity.
Appendix E Numerical Validations for Semi-Norm Contractions
In this section, we provide numerical examples that directly verify the one-step strict contraction across the settings studied. These results empirically support the key structural claims used by our analysis.
E.1 Evaluations of Lemma 4.1
Lemma A.2 is the technical backbone for Lemma 4.1, as Lemma A.2 constructs the fixed-kernel semi-norm and provides the one-step contraction for a given $P$. Therefore, we perform numerical evaluations on Lemma A.2 to demonstrate the one-step contraction property for Lemma 4.1.
For a kernel $P$ with stationary distribution $d$, we follow the steps in Appendix A.1 and construct $\lVert\cdot\rVert_{\mathsf{P}}$ in (41) with
$\alpha=\min\{0.99,\tfrac{1+\rho(Q)}{2}\},\qquad\epsilon=0.25(1-\alpha),$ (192)
where $Q=P-\mathbf{e}\,d^{\top}$. Then the one-step contraction factor is $\beta=\alpha+\epsilon<1$.
To generate ergodic matrices with dimension $n$, let $I_{n}$ be the identity matrix and $S_{n}$ be the cyclic shift matrix defined by $S_{n}e_{i}=e_{i+1\bmod n}$. We provide the following four examples:
  • $P_{1}=0.5\,I_{5}+0.5\,S_{5}$
  • $P_{2}=0.6\,I_{6}+0.4\,S_{6}$
  • $P_{3}=0.55\,I_{7}+0.45\,S_{7}$
  • $P_{4}=0.6\,I_{8}+0.3\,S_{8}+0.1\,S_{8}^{2}$
We generate $1000$ random unit vectors $x$ and compute each ratio $\frac{\lVert P_{i}x\rVert_{\mathsf{P}}}{\lVert x\rVert_{\mathsf{P}}}$. The empirical results are summarized below.
matrix $n$ max span ratio $\rho(Q)$ $\alpha$ $\epsilon$ $\beta$ $\mathrm{ratio}_{\min}$ $\mathrm{ratio}_{\mathrm{median}}$ $\mathrm{ratio}_{\mathrm{p90}}$ $\mathrm{ratio}_{\max}$
P1 5 1 0.8090 0.9045 0.0239 0.9284 0.3824 0.7950 0.8077 0.8090
P2 6 1 0.8718 0.9359 0.0160 0.9519 0.5197 0.8510 0.8687 0.8718
P3 7 1 0.9020 0.9510 0.0122 0.9633 0.5861 0.8799 0.8976 0.9014
P4 8 1 0.8700 0.9350 0.0162 0.9513 0.4855 0.8226 0.8604 0.8685
Table 1: Empirical one-step contraction ratios for ergodic kernels using the fixed-kernel semi-norm $\lVert\cdot\rVert_{\mathsf{P}}$.
E.2 Evaluations of Theorem 4.2
Lemma A.8 is the key step for Theorem 4.2, as Lemma A.8 proves a uniform one-step contraction across all $P$ in the uncertainty set. We therefore perform numerical evaluations on Lemma A.8 to demonstrate the one-step contraction property for Theorem 4.2 under contamination, total variation (TV), and Wasserstein-1 uncertainty. We select the same $P_{1}$, $P_{2}$, $P_{3}$, and $P_{4}$ in Appendix E.1 as four examples of the nominal model.
To numerically approximate $\lVert\cdot\rVert_{\mathcal{P}}$ defined in (82), we approximate $\lVert\cdot\rVert_{\mathcal{P}}$ by (i) discretizing the uncertainty set and (ii) using a finite product to approximate the extremal norm. First, we sample a family $\{P^{(i)}\}_{i=1}^{m}\subset\mathcal{P}$ of size $m$ and form
$Q_{i}\;:=\;P^{(i)}-\mathbf{e}\,d_{P^{(i)}}^{\top},\qquad\hat{r}\;=\;\max_{i}\rho(Q_{i}).$
We set $\alpha=\min\{0.99,(1+\hat{r})/2\}$ and choose $\epsilon\in(0,1-\alpha)$. To approximate the extremal norm, we build a library of scaled products of the $Q_{i}$’s up to maximum length $K$: for each $k=0,1,\dots,K$ we draw products $M_{k,j}=Q_{i_{k}}\cdots Q_{i_{1}}$; the number of such draws at each $k$ is the “products per length” (denoted samples_per_k in the tables). This defines the surrogate
$\lVert z\rVert_{\mathrm{ext}}^{(K)}=\max_{0\leq k\leq K,\,j}\ \alpha^{-k}\,\lVert M_{k,j}z\rVert_{2}.$
We then set
$\lVert x\rVert_{\mathcal{P}}^{(K)}\;=\;\max_{i=1,\dots,m}\ \lVert Q_{i}x\rVert_{\mathrm{ext}}^{(K)}\;+\;\varepsilon\,\min_{c\in\mathcal{C}}\ \lVert x-c\,\mathbf{e}\rVert_{\mathrm{ext}}^{(K)}.$
We generate $50$ random unit vectors $x$ for each sampled uncertainty matrix in $\{P^{(i)}\}_{i=1}^{m}\subset\mathcal{P}$, and compute the ratios $\frac{\lVert P^{(i)}x\rVert_{\mathcal{P}}}{\lVert x\rVert_{\mathcal{P}}}$. The empirical results of the uncertainty sets studied in our settings are summarized below.
nominal $\tilde{\mathsf{P}}$ $n$ $\delta$ $m$ $K$ samples_per_k max span ratio $\hat{r}$ $\alpha$ $\epsilon$ $\gamma$ $\mathrm{ratio}_{\min}$ $\mathrm{ratio}_{\mathrm{median}}$ $\mathrm{ratio}_{\mathrm{p90}}$ $\mathrm{ratio}_{\max}$
P1 5 0.15 30 3 25 1 0.8138 0.9069 0.0233 0.9302 0.2210 0.6396 0.8057 0.8134
P2 6 0.15 30 3 25 1 0.8807 0.9403 0.0149 0.9553 0.2957 0.6581 0.8630 0.8785
P3 7 0.15 30 3 25 1 0.9067 0.9534 0.0117 0.9650 0.3217 0.6683 0.8700 0.8901
P4 8 0.15 30 3 25 1 0.8812 0.9406 0.0148 0.9555 0.3739 0.5964 0.8207 0.8624
Table 2: Empirical one-step contraction ratios under contamination uncertainty.
nominal $\tilde{\mathsf{P}}$ $n$ $\delta$ $m$ $K$ samples_per_k max span ratio $\hat{r}$ $\alpha$ $\epsilon$ $\gamma$ $\mathrm{ratio}_{\min}$ $\mathrm{ratio}_{\mathrm{median}}$ $\mathrm{ratio}_{\mathrm{p90}}$ $\mathrm{ratio}_{\max}$
P1 5 0.15 30 3 25 1 0.8280 0.9140 0.0215 0.9355 0.3239 0.7609 0.8239 0.8280
P2 6 0.15 30 3 25 1 0.9013 0.9507 0.0123 0.9630 0.4021 0.7904 0.8862 0.9006
P3 7 0.15 30 3 25 1 0.9175 0.9588 0.0103 0.9691 0.4457 0.7918 0.8962 0.9162
P4 8 0.15 30 3 25 1 0.8805 0.9403 0.0149 0.9552 0.4497 0.7503 0.8449 0.8739
Table 3: Empirical one-step contraction ratios under total variation (TV) uncertainty.
nominal $\tilde{\mathsf{P}}$ $n$ $\delta$ $m$ $K$ samples_per_k max span ratio $\hat{r}$ $\alpha$ $\epsilon$ $\gamma$ $\mathrm{ratio}_{\min}$ $\mathrm{ratio}_{\mathrm{median}}$ $\mathrm{ratio}_{\mathrm{p90}}$ $\mathrm{ratio}_{\max}$
P1 5 0.15 30 3 25 1 0.8184 0.9092 0.0227 0.9319 0.3698 0.7569 0.8141 0.8188
P2 6 0.15 30 3 25 1 0.8900 0.9450 0.0138 0.9587 0.4257 0.7889 0.8774 0.8894
P3 7 0.15 30 3 25 1 0.9110 0.9555 0.0111 0.9666 0.4262 0.7818 0.8827 0.9080
P4 8 0.15 30 3 25 1 0.8758 0.9379 0.0155 0.9534 0.4413 0.7224 0.8518 0.8689
Table 4: Empirical one-step contraction ratios under Wasserstein-1 uncertainty.
E.3 Interpretations
[…]
Appendix F Some Auxiliary Lemmas for the Proofs
Lemma F.1 (Theorem IV in [2]).
@@ Appendix F Some Auxiliary Lemmas for the Proofs
Let $\mathcal{Q}$ be a bounded set of square matrix such that $\rho(Q)<\infty$ for all $Q\in\mathcal{Q}$ where $\rho(\cdot)$ denotes the spectral radius. Then the joint spectral radius of $\mathcal{Q}$ can be defined as
$\hat{\rho}(\mathcal{Q})\coloneqq\lim_{k\rightarrow\infty}\sup_{Q_{i}\in\mathcal{Q}}\rho(Q_{k}\ldots Q_{1})^{\frac{1}{k}}=\lim_{k\rightarrow\infty}\sup_{Q_{i}\in\mathcal{Q}}\|Q_{k}\ldots Q_{1}\|^{\frac{1}{k}}$ (193)
where $\|\cdot\|$ is an arbitrary norm.
Lemma F.2 (Lemma 6 in [42]).
@@
Under the setup and notation in Appendix B.1.1, if assuming the noise has bounded variance of $\mathbb{E}[\,\|w^{t}\|_{\mathcal{N},\overline{E}}^{2}|\mathcal{F}^{t}]\leq A+B\|x^{t}-x^{*}\|_{\mathcal{N},\overline{E}}^{2}$, we have
@@
Lemma F.4 (Ergodic case of Lemma 9 in [33]).
For any average-reward MDP with stationary policy $\pi$ and the mixing time defined as